two hamburgers and one bagel cost $4.82.
At the same price, one hamburger and
two bagels cost $3.70. How much does
each hamburger and each bagel cost?

Found 2 solutions by richwmiller, brucewill:
Answer by richwmiller(17219)

2h+1b=4.82
1h+2b=3.70
h=1.98 and b=.86

Answer by brucewill(101)

Let x = number of hamburgers
Let y = number of bagels
2x + y = 482
x + 2y = 370
--- multiply second equation by 2
2x + y = 482
2x + 4y = 740
--- subtract second equation from first
-3y = -258
---
y = 86; substitute into first equation:
---
2x + 86 = 482
2x = 396
x = 198
---
Proof: 2*(198) + 86 = 396 + 86 = 482
198 + 2(86) = 198 + 172 = 370